Suppose that the reserve requirement in an economy is 0.10 and the total amount of excess reserves is 100 billion. Whac is the maximum possible total change in the money supply
Answer #1

We know money multiplier=1/r=1/0.1=10

excess reserve is 100 billion

then maximum possible total change in money supply=10*100=1000billion